In an automobile race the cars are started with a starter car traveling at 31.5 m/sec. If

the lead car has an average acceleration of 4.3 m/sec2 during the first 3 seconds, how
fast was it traveling at the end of 3 seconds and how far did it go?
1
Expert's answer
2011-10-14T09:19:28-0400
V=31.5m/sec
a=4.3m/sec^2
v(after 3sec)=V+at=31.5+4.3*3=44.4
s(after 3sec)=Vt+at^2/2=31.5*3+0.5*4.3*9=94.5+19.35=113.85metres